# SMR: State Memory Replay for Long Sequence Modeling

### v1

@@ Abstract
Biqing Qi1,2,, Junqi Gao3,thanks: Equal contributions., Kaiyan Zhang2, Dong Li3, Jianxing Liu1, Ligang Wu1,thanks: Corresponding authors: Bowen Zhou and Ligang Wu., Bowen Zhou2, 1 Department of Control Science and Engineering, Harbin Institute of Technology, 2 Department of Electronic Engineering, Tsinghua University, 3 School of Mathematics, Harbin Institute of Technology, {qibiqing7,gjunqi97,arvinlee826}@gmail.com, zhang-ky22@mails.tsinghua.edu.cn, {jx.liu,ligangwu}@hit.edu.cn, zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Long sequence modeling has attracted extensive interest due to its broad prospects in natural language processing (Beltagy et al., 2020; Brown et al., 2020; Ouyang et al., 2022). The mainstream architectures for sequence modeling mainly focus on attention-based Transformers (Vaswani et al., 2017). However, the quadratic complexity of softmax attention brings a computational bottleneck (Choromanski et al., 2020; Wang et al., 2020; Beltagy et al., 2020), which makes attention-based architectures inefficient for handling long sequences. Although the introduction of linear attention (Wang et al., 2020) reduces the computational complexity, it cannot well approximate the performance of the vanilla Transformer. More importantly, purely attention-based architectures cannot capture long-range dependencies well. On the other hand, state space model (SSM)-based architectures (Gu et al., 2021a; Gupta et al., 2022) show superior performance on the Long Range Arena (LRA) (Tay et al., 2021) benchmark for long sequence modeling due to their linear computational complexity and excellent long-range dependency capturing ability. Existing SSM-based model architectures, such as S5 (Smith et al., 2023) and S6 (Gu and Dao, 2023), primarily rely on recursive structures to tackle the varying sampling step issue. S5 introduced learnable step sizes for each step to improve the Sampling Step Adaptation (SSA) capability of SSM. S6 (Mamba) introduced data-dependent parameter settings, which makes the state propagation of the SSM model more flexible. However, this restricts its inference computation to parallel scanning instead of the original efficient convolution mode, significantly hampering training efficiency and imposing a heavier inference burden when handling long inputs at once.
@@ 3.2 Theoretical Understanding of NSS
Figure 2: An illustrative instance of the NSS issue in S4 is presented here. $\text{States}_{raw}$ and $\text{States}_{Pert}$ denote the value of latent states of the model when applied to the original fitting data and the data subject to sampled perturbations, respectively.
@@ 3.4 Empirical Validation of SMR for SSA
Figure 4: Comparative validation of S4 incorporated with SMR (S4+SMR) on the aforementioned examples. The pair of figures displays the prediction outcomes of S4+SMR for the perturbed input $u^{\prime}$ (left) and the latent states of the model when provided with inputs $u$ and $u^{\prime}$ (right).
@@ 6 Limitations
However, the current study is preliminary. Future work will explore utilizing more advanced control theory to enhance the proposed mechanism. Additionally, we will investigate the performance of the SMR mechanism under different task types and data distributions.

### v2

@@ Abstract
Biqing Qi1,2,4,, Junqi Gao3,thanks: Equal contributions., Kaiyan Zhang2, Dong Li3, Jianxing Liu1, Ligang Wu1,thanks: Corresponding authors: Bowen Zhou and Ligang Wu., Bowen Zhou2, 1 Department of Control Science and Engineering, Harbin Institute of Technology, 2 Department of Electronic Engineering, Tsinghua University, 3 School of Mathematics, Harbin Institute of Technology, 4 Frontis.AI, Beijing {qibiqing7,gjunqi97,arvinlee826}@gmail.com, zhang-ky22@mails.tsinghua.edu.cn, {jx.liu,ligangwu}@hit.edu.cn, zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Long sequence modeling has attracted extensive interest due to its broad prospects in natural language processing (Beltagy et al., 2020; Brown et al., 2020; Ouyang et al., 2022). The mainstream architectures for sequence modeling mainly focus on attention-based Transformers (Vaswani et al., 2017). However, the quadratic complexity of softmax attention brings a computational bottleneck (Choromanski et al., 2020; Wang et al., 2020; Beltagy et al., 2020), which makes attention-based architectures inefficient for handling long sequences. Although the introduction of linear attention (Wang et al., 2020) reduces the computational complexity, it cannot well approximate the performance of the vanilla Transformer. More importantly, purely attention-based architectures cannot capture long-range dependencies well. On the other hand, state space model (SSM)-based architectures (Gu et al., 2021a; Gupta et al., 2022) show superior performance on the Long Range Arena (LRA) (Tay et al., 2021) benchmark for long sequence modeling due to their linear computational complexity and excellent long-range dependency capturing ability.
Existing SSM-based model architectures, such as S5 (Smith et al., 2023) and S6 (Gu and Dao, 2023), primarily rely on recursive structures to tackle the varying sampling step issue. S5 introduced learnable step sizes for each step to improve the Sampling Step Adaptation (SSA) capability of SSM. S6 (Mamba) introduced data-dependent parameter settings, which makes the state propagation of the SSM model more flexible. However, this restricts its inference computation to parallel scanning instead of the original efficient convolution mode, significantly hampering training efficiency and imposing a heavier inference burden when handling long inputs at once.
@@ 3.2 Theoretical Understanding of NSS
Figure 2: An illustrative instance of the NSS issue in S4 is presented here.
@@ 3.4 Empirical Validation of SMR for SSA
Figure 4: Comparative results of S4 incorporated with SMR (S4+SMR) on the aforementioned examples. The pair of figures displays the prediction outcomes of S4+SMR for the perturbed input $u^{\prime}$ (left) and the latent states when provided with inputs $u$ and $u^{\prime}$ (right).
@@ 6 Limitations
However, the current study is preliminary. In the future, we can extend this technology to interactive learning frameworks Qi et al. (2024a), explore continual SSM frameworks Qi et al. (2024b), and design more robust and secure models Qi et al. (2024d); Gao et al. (2023); Qi et al. (2024c), applying them to scenarios such as knowledge discovery Qi et al. (2023).
@@ 7 Ethics Statement
8 acknowledgement
This work was supported in part by the National Science and Technology Major Project (No. 20232D0121403). We extend our gratitude to the anonymous reviewers for their insightful feedback, which has greatly contributed to the improvement of this paper.